A car is travelling at a constant speed of 350m/s. Passes a trooper hidden behind a Billboard one second of the car passes the Billboard, the trooper sets out from the Billboard to catch it, acelerating at a constant speed of 30m/s².how long does it take her to take the car?
1
Expert's answer
2020-02-17T09:26:25-0500
"d=v(t+1)"


"d=0.5at^2"

Thus,


"v(t+1)=0.5at^2"


"350(t+1)=0.5(30)t^2"

"t=24\ s"